An open pipe 30cm long and a closed pipe 23cm long, both of same diameter , are each sounding their first overtone are in unison . The end correction of these pipes is

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Suppose v Is the velocity of light and n is Me frequency of the note.When tho open lalos io wunging its first overtone. the length of the pipe plus end correction = 1 = =30 1.2cSince there are Wo end cone-When Me closed pipe is sound- ing its first overtone 4' = c=23+c or,n = 23+c •00 From 01 en (0) 23 +0 = 1=00 1.28)92+4c = 90+ 8c c = 1 cm